# Enhancing Accuracy in Differentially Private Distributed Optimization Through Sensitivity Reduction

### v1

@@ 9 Proof of Basic Lemmas
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{x}(k)\|^{2}|\mathcal{M}_{k}\right]$ (76)
@@
$\quad+\mathbb{E}\left[\|\bm{d}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},\bm{b}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{b}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{b}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{c}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right].$
@@
$\quad W(\mathbf{x}(k)-\mathbf{x}(k-1))$ (77)
@@
$\quad+\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k).$
@@
$\quad\mathbb{E}\left[\|\bm{a}_{3}\|^{2}|\mathcal{M}_{k}\right]$ (78)
@@
$\quad+\alpha_{k+1}^{2}\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+2q_{1}\mathbb{E}\Big{[}\Big{\langle}(W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T})(\mathbf{x}(k)-\mathbf{x}(k-1)),$
$\qquad-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2q_{1}^{2}\mathbb{E}\Big{[}\Big{\langle}(W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T})\mathbf{x}(k),\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2q_{1}\mathbb{E}\Big{[}\Big{\langle}-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
@@
$\quad+\sigma\Big{(}\frac{1-\sigma^{2}}{4\sigma}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+\frac{4\sigma}{1-\sigma^{2}}\mathbb{E}\Big{[}\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\Big{]}\Big{)}$
$\quad+2np(7+2L\alpha_{k}+2L\alpha_{k+1})\nu_{k}^{2}$
@@
$\quad+\frac{1+3\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\Big{[}\Big{\|}\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+2np(7+2L\alpha_{k}+2L\alpha_{k+1})\nu_{k}^{2},$
@@
$\quad\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]$ (79)
@@
$\quad+2npL^{2}\alpha_{k+1}^{2}\nu_{k+1}^{2}+2npL^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}\nu_{k}^{2}.$
@@
$\quad+\frac{\alpha_{k+1}^{2}}{\alpha_{k}^{2}}\mathbb{E}\left[\left\|W\bm{\xi}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\leq 18np\nu_{k+1}^{2}+2np\nu_{k}^{2}.$
@@
$\quad\mathbb{E}\left[2\langle\bm{a}_{3},\bm{b}_{3}\rangle|\mathcal{M}_{k}\right]$ (82)
@@
$\quad+2\mathbb{E}\left[\left\langle-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\bm{b}_{3}\right\rangle|\mathcal{M}_{k}\right]$
$\quad+2q_{1}\mathbb{E}\Big{[}\Big{\langle}\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k),(\tilde{W}(k)-I)\mathbf{x}(k)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
@@
$\quad+\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\Big{[}\Big{\|}\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
$\quad+8np(3+L\alpha_{k})\nu_{k}^{2}$
@@
$\quad+\mathbb{E}\Big{[}\Big{\|}\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}+8np(3+L\alpha_{k})\nu_{k}^{2},$
@@
$\quad\mathbb{E}\left[2\langle\bm{a}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]$ (83)
@@
$\quad+2\mathbb{E}\left[\left\langle\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\bm{c}_{3}\right\rangle|\mathcal{M}_{k}\right]$
$\quad+2q_{1}\alpha_{k+1}\mathbb{E}\Big{[}\Big{\langle}\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k),\nabla F(\mathbf{z}(k))-\nabla F(\mathbf{z}(k+1))\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
@@
$\quad+\mathbb{E}\left[\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2}$
@@
$\quad+\mathbb{E}\left[\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2},$
@@
$\quad\mathbb{E}\left[2\langle\bm{b}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$ (86)
@@
$\quad\mathbb{E}\left[2\langle-\bm{c}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$ (87)
@@
$\qquad\left((1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\right)\bm{\xi}(k+1)\rangle|\mathcal{M}_{k}]$
$\quad+2\mathbb{E}\left[\left\langle\alpha_{k+1}\nabla F(\mathbf{z}(k+1)),\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\quad+2\mathbb{E}\left[\left\langle\alpha_{k+1}\nabla F(\mathbf{z}(k)),-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 12npL\alpha_{k+1}\nu_{k+1}^{2}+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2}.$
@@
$\quad\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$ (88)
@@
$\quad+2\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k+1))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
@@
$\quad+\frac{4L^{2}}{n}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
@@
$\quad+\frac{4L^{2}}{n}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]+12pL^{2}\nu_{k+1}^{2}$
$\quad+4pL^{2}(7+2L\alpha_{k})\nu_{k}^{2},$
@@
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{x}(k)\|^{2}|\mathcal{M}_{k}\right]$ (89)
@@
$\quad+4np(19+6L\alpha_{k}+5L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1})\nu_{k}^{2}$
$\quad+\frac{3+\sigma^{2}}{1-\sigma^{2}}\Big{(}\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+n\alpha_{k+1}^{2}\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}\Big{)}$
@@
$\quad+\frac{4n(3+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{(3+\sigma^{2})(\|W-I\|^{2}+4L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+4np\Big{(}19+6L\alpha_{k}+5L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}$

### v2

@@ 9 Proof of Basic Lemmas
$\qquad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{x}(k)\|^{2}|\mathcal{M}_{k}\right]$ (76)
@@
$\qquad+\mathbb{E}\left[\|\bm{d}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},\bm{b}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]$
$\qquad+\mathbb{E}\left[2\langle\bm{b}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$
$\qquad+\mathbb{E}\left[2\langle\bm{b}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{c}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right].$
@@
$\qquad W(\mathbf{x}(k)-\mathbf{x}(k-1))$ (77)
@@
$\qquad+\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k).$
@@
$\qquad\mathbb{E}\left[\|\bm{a}_{3}\|^{2}|\mathcal{M}_{k}\right]$ (78)
@@
$\qquad+\alpha_{k+1}^{2}\mathbb{E}\Big[\Big\|\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big\|^{2}|\mathcal{M}_{k}\Big]$
$\qquad+2q_{1}\mathbb{E}\Big[\Big\langle(W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T})(\mathbf{x}(k)-\mathbf{x}(k-1)),$
$\hskip 17.00024pt-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big\rangle|\mathcal{M}_{k}\Big]$
$\qquad+2q_{1}^{2}\mathbb{E}\Big[\Big\langle(W-\frac{1}{n}\mathbf{1}\mathbf{1}^{T})\mathbf{x}(k),\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k)\Big\rangle|\mathcal{M}_{k}\Big]$
$\qquad+2q_{1}\mathbb{E}\Big[\Big\langle-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k)\Big\rangle|\mathcal{M}_{k}\Big]$
@@
$\qquad+\sigma\Big(\frac{1-\sigma^{2}}{4\sigma}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\hskip 17.00024pt+\frac{4\sigma}{1-\sigma^{2}}\mathbb{E}\Big[\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\Big]\Big)$
$\qquad+2np(7+2L\alpha_{k}+2L\alpha_{k+1})\nu_{k}^{2}$
@@
$\qquad+\frac{1+3\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\Big[\Big\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big\|^{2}|\mathcal{M}_{k}\Big]$
$\qquad+2np(7+2L\alpha_{k}+2L\alpha_{k+1})\nu_{k}^{2},$
@@
$\qquad\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]$ (79)
@@
$\qquad+2npL^{2}\alpha_{k+1}^{2}\nu_{k+1}^{2}+2npL^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}\nu_{k}^{2}.$
@@
$\qquad+\frac{\alpha_{k+1}^{2}}{\alpha_{k}^{2}}\mathbb{E}\left[\left\|W\bm{\xi}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
[…]
@@
$\qquad\mathbb{E}\left[2\langle\bm{a}_{3},\bm{b}_{3}\rangle|\mathcal{M}_{k}\right]$ (82)
@@
$\qquad+2\mathbb{E}\left[\left\langle-\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\bm{b}_{3}\right\rangle|\mathcal{M}_{k}\right]$
$\qquad+2q_{1}\mathbb{E}\Big[\Big\langle\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k),(\tilde{W}(k)-I)\mathbf{x}(k)\Big\rangle|\mathcal{M}_{k}\Big]$
@@
$\qquad+\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\Big[\Big\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big\|^{2}|\mathcal{M}_{k}\Big]$
$\qquad+8np(3+L\alpha_{k})\nu_{k}^{2}$
@@
$\qquad+\mathbb{E}\Big[\Big\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big\|^{2}|\mathcal{M}_{k}\Big]+8np(3+L\alpha_{k})\nu_{k}^{2},$
@@
$\qquad\mathbb{E}\left[2\langle\bm{a}_{3},-\bm{c}_{3}\rangle|\mathcal{M}_{k}\right]$ (83)
@@
$\qquad+2\mathbb{E}\left[\left\langle\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k)),\bm{c}_{3}\right\rangle|\mathcal{M}_{k}\right]$
$\qquad+2q_{1}\alpha_{k+1}\mathbb{E}\Big[\Big\langle\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\bm{\xi}(k),\nabla F(\mathbf{z}(k))-\nabla F(\mathbf{z}(k+1))\Big\rangle|\mathcal{M}_{k}\Big]$
@@
$\qquad+\mathbb{E}\left[\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2}$
@@
$\qquad+\mathbb{E}\left[\left\|\frac{\alpha_{k+1}}{n}\mathbf{1}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\right\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2},$
@@
$\qquad\mathbb{E}\left[2\langle\bm{b}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$ (86)
@@
$\qquad\mathbb{E}\left[2\langle-\bm{c}_{3},\bm{d}_{3}\rangle|\mathcal{M}_{k}\right]$ (87)
@@
$\hskip 18.49988pt\left((1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\right)\bm{\xi}(k+1)\rangle|\mathcal{M}_{k}]$
$\qquad+2\mathbb{E}\left[\left\langle\alpha_{k+1}\nabla F(\mathbf{z}(k+1)),\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\qquad+2\mathbb{E}\left[\left\langle\alpha_{k+1}\nabla F(\mathbf{z}(k)),-\frac{\alpha_{k+1}}{\alpha_{k}}W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
[…]
@@
$\qquad\mathbb{E}\Big[\Big\|\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big\|^{2}|\mathcal{M}_{k}\Big]$ (88)
@@
$\qquad+2\mathbb{E}\Big[\Big\|\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k+1))\Big\|^{2}|\mathcal{M}_{k}\Big]$
@@
$\qquad+\frac{4L^{2}}{n}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
@@
[…]
$\qquad+4pL^{2}(7+2L\alpha_{k})\nu_{k}^{2},$
@@
$\qquad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{x}(k)\|^{2}|\mathcal{M}_{k}\right]$ (89)
@@
[…]
$\qquad+\frac{3+\sigma^{2}}{1-\sigma^{2}}\Big(\mathbb{E}\left[\|\bm{b}_{3}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[\|\bm{c}_{3}\|^{2}|\mathcal{M}_{k}\right]$
$\hskip 17.00024pt+n\alpha_{k+1}^{2}\mathbb{E}\Big[\Big\|\frac{1}{n}\mathbf{1}^{T}\nabla F(\mathbf{z}(k))\Big\|^{2}|\mathcal{M}_{k}\Big]\Big)$
@@
$\qquad+\frac{4n(3+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+\frac{(3+\sigma^{2})(\|W-I\|^{2}+4L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+2np\Big(9+6L\alpha_{k+1}+\frac{28L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big)\nu_{k+1}^{2}$
[…]
$\hskip 17.00024pt+\frac{6L^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big)\nu_{k}^{2}.$